# DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

### v2

@@ 3.2 Adaptive Modality-Specific Mixture-of-Experts
The adaptive modality expert selection strategy is primarily designed for an original textual LLM, dynamically distinguishing between audio experts and text experts based on modality load. The goal is to ensure minimal impairment to the language capabilities of the original LLM. Specifically, as shown in Algorithm 1, an initial MoE model that has undergone stage 1 modality alignment training is used. For unimodal input data, the token load rates of all experts for both modalities are calculated, as illustrated in Figure 4. According to step 1 in Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts. As shown in Figure 4, for example, in the first layer, the expert with index 5 (starting from 0) is selected as an audio expert because its load rate for the audio modality is significantly higher than that for the text modality. This adaptive selection approach fully utilizes the text experts with low load in the original MoE LLM, while also ensuring that the selected experts are suitable for processing the audio modality. In this way, the strategy minimizes the impact on the original LLM while fitting the audio modality as much as possible.
Algorithm 1 Adaptive Modality Expert Partitioning and Training
0:  Data with audio input and audio-text output $D^{A}$, Data with text input and audio-text output $D^{T}$, The MoE LLM model $M$ after the first stage of modality alignment.
1:  Initialize $k$ = Number of audio experts, $N$ = Total number of experts.
2:  Step 1. Adaptive partitioning of modality experts
3:   Audio modality token load rate $E^{A}$ = $M(D^{A})$   # Inference with model $M$ on dataset $D^{A}$.
4:   Text modality token load rate $E^{T}$ = $M(D^{T})$   # Inference with model $M$ on dataset $D^{T}$.
5:   _$\_,AE$ = $topk(E^{A}*(1-E^{T}),k)$   # Compute the index $AE$ of the selected audio expert.
6:   _$\_,TE$ = $topk(E^{T}*(1-E^{A}),N-k)$   # Compute the index $TE$ of the selected text expert.
7:  Step 2. Expert specialization training for single modality
8:   $M[AE]$ = $M[AE](D^{A})$   # The audio expert undergoes specialized training on $D^{A}$.
9:   $M[TE]$ = $M[TE](D^{T})$   # The text expert undergoes specialized training on $D^{T}$.
10:  Stpe 3. Joint training of multi-modality experts
11:   $M^{\prime}$ = $M(D^{A}\cup D^{T})$
12:  return $M^{\prime}$, the MoE SLM model after modality expert partitioning and training.
@@ 3.3 Training Strategy for Modality Experts
Modality Aligment During the modality alignment stage, audio ASR data is used to align the semantic spaces of speech and text. In this stage, the main focus is on training the Adapter that connects the Audio Encoder and the LLM. Through modality alignment training, it is ensured that speech can be understood by the LLM in the same way as the meaning expressed by text.
@@ 3.4 Audio Generation with Reinforcement Learning
To broaden the coverage of speakers and domains, it is inevitable that the pre-training data contains label noise and pronunciation errors, which can lead to hallucinations in the model. We observed that the quality of generated response audio is not consistent under different sampling parameters. To ensure more stable quality in response audio generation, we introduce a reinforcement learning stage to improve the stability of speech generation. Specifically, we use the text from the LibriSpeech panayotov2015librispeech set and have the model paraphrase it to obtain response audio and corresponding text. The response audio is then transcribed into recognized text using Whisper-large, and the word error rate (WER) is calculated as the reward score. All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) rafailov2023direct training as follows:
@@
where $\theta$ is the policy model and $ref$ is the reference model. $x$ denotes the text input of the audio to be synthesized, $y_{w}^{A}$ denotes a good speech sequence, and $y_{l}^{A}$ denotes a bad speech sequence.
@@ 4.1 Experiment Settings
Models The overall architecture of the DeepTalk model is shown in Figure 2. We use the pre-trained DeepSeek-V2-Lite as the initial MoE-based LLM backbone, Whisper-medium as the audio encoder, and an MLP-based adapter to connect the audio encoder and the LLM. DeepTalk contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts. A sparse routing mechanism is adopted, where each token is processed by 6 experts. The model consists of 27 layers in total: the first layer is a dense layer, followed by 26 MoE layers. Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention dao2023flashattention with 16 attention heads. In stage 1, the learning rate is set to 2e-5. Since the original LLM has not undergone any pre-training on audio modalities, the learning rate is set to 1e-4 in stage 2.1 to accelerate convergence, 2e-5 in stage 2.2, and 5e-5 in stage 3. We use the AdamW loshchilov2017decoupled optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95. The final evaluation is conducted on LLM benchmarks, SQA, ASR, and TTS tasks.
Datasets In stage 1, WenetSpeech zhang2022wenetspeech is used for modality alignment. In stage 2.1, AudioQA-1M gao2025lucy is used for specialized training of audio experts. In stage 2.2, inhouse-60K, databricks-dolly-15k dolly2023introducing , MathInstruct yue2023mammoth (262K) and camel-ai-math li2023camel (50K) are used for specialized training of text experts. In stage 3, AudioQA-1M is used for joint training of modality experts. For the RL stage, we sampled 28K text entries from LibriSpeech panayotov2015librispeech to construct the audio preference pair dataset.
@@ 4.2 Evaluations
Evaluation on Text to Text
Figure 4: The modality load of the experts in the first layers.
The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/. As shown in Table 3, modular MLLMs cause minimal damage to the language abilities of the original text LLMs, primarily because the LLM is only responsible for generating text tokens. However, native MLLMs significantly impair the language capabilities of the original text LLMs, with an average relative loss of over 20% compared to their backbone LLMs. The DeepTalk model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
@@
Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo chen2025minmo 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio huang2025step 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni wang2024freeze 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi defossez2024moshi 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice zeng2024glm 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY gao2025lucy 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend zhong2024moextend 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepTalk 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
The evaluation of speech-to-text is divided into two tasks: Spoken QA and ASR. As shown in the S $\rightarrow$ T column of Table 2, thanks to the preservation of the original LLM’s language capabilities by modular MLLMs, these models generally achieve leading performance. In contrast, native MLLMs cause significant degradation to the original LLM, resulting in increased hallucinations and thus lower scores. However, DeepTalk effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent. As a result, DeepTalk leads among native MLLMs. The ASR capabilities of the models are shown in Table 4, where DeepTalk, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
@@
Model #A-Params WenetSpeech AIShell LibriSpeech
test-net test-meeting test test-clean test-other
Qwen2-Audio-base chu2024qwen2 7B 7.64 8.40 1.52 1.74 4.04
Baichuan-Audio-base li2025baichuan 7B 10.13 13.28 1.93 3.02 6.04
VITA fu2025vita 7B 12.15 16.53 - 8.14 18.41
Step-Audio-chat huang2025step 130B 9.47 10.83 2.14 3.19 10.67
Qwen2.5-Omni xu2025qwen2 7B 6.04 7.71 1.13 2.37 4.21
GLM-4-Voice zeng2024glm 9B - - 3.02 2.10 4.90
DeepTalk 2.4B 8.98 10.67 2.42 3.25 8.21
Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepTalk, where "topk" denotes the sampling parameter for audio tokens. Since the speech dialogue data used by DeepTalk contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English. After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in Recall@1, thereby making high-quality audio generation more stable. However, for more difficult tasks (test-hard), the supervised fine-tuning (SFT) model itself may not generate the correct result within the entire search space, so applying DPO brings little to no benefit.
Evaluation on Speech to Speech The evaluation of speech-to-speech is shown in the S $\rightarrow$ S column of Table 2. MoExtend refers to the results obtained by extending the original Deepseek-V2-Lite model with additional experts, adding six audio experts. However, since expanding the number of experts alters the distribution of the router, the performance is relatively suboptimal. DeepTalk, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts. This allows the audio experts to specialize while retaining the capabilities of the original text experts, thus achieving better performance among native MLLMs.
Latency We deployed DeepTalk on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module. The reported numbers are averaged over more than 10 samples. With a good network connection, the end-to-end latency of the entire system in half-duplex mode is 0.4362 seconds. The time cost for generating the first audio chunk is 0.3417 seconds. By subtracting the first chunk cost from the total latency, we estimate the network transmission cost to be 0.1062 seconds. The time cost for decoding a single token is 0.0205 seconds. All time measurements were conducted on a GPU with 80GB of memory.
@@ 5 Conclusion
To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepTalk, a MoE framework for adaptive modality expert learning. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs a performance loss lower than peer native MLLMs and on par with modular ones. Meanwhile, the end-to-end dialogue latency remains within 0.5 seconds, ensuring a seamless and intelligent speech interaction experience. To the best of our knowledge, this is the first MoE-based native multimodal large speech interaction model, which opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
Appendix A Dataset Details
@@ Appendix B Experimental Setup Details
Table 6: Hyperparameters used in each stage training on DeepTalk.
@@
Appendix C Ablation Study
C.1 Audio Experts Number
@@
C.2 The Modality Load of Different Layer Experts
@@
The expert load across different layers is illustrated in Figures 5 and 6, representing the modality load of an intermediate MoE layer and the final MoE layer, respectively. It is evident that there are significant differences in modality load across different layers. This indicates that the indices of audio experts in each layer are not the same.
C.3 Modality-Specific MoE

### v3

@@ 3.2 Adaptive Modality-Specific Mixture-of-Experts
The adaptive modality expert selection strategy is primarily designed for an original textual LLM, dynamically distinguishing between audio experts and text experts based on modality load. The goal is to ensure minimal impairment to the language capabilities of the original LLM. Specifically, as shown in Algorithm 1, an initial MoE model that has undergone stage 1 modality alignment training is used. For unimodal input data, the token load rates of all experts for both modalities are calculated. According to Algorithm 1, those experts with low text token load but high audio token load are selected as audio experts, while the others are designated as text experts. This adaptive selection approach fully utilizes the text experts with low load in the original MoE LLM, while also ensuring that the selected experts are suitable for processing the audio modality. In this way, the strategy minimizes the impact on the original LLM while fitting the audio modality as much as possible.
Algorithm 1 Adaptive Modality Expert Partitioning
1:
  • Multimodal dataset with audio input: $\mathcal{D}^{A}=\{(x_{i}^{A},y_{i}^{A,T})\}_{i=1}^{n}$
  • Multimodal dataset with text input: $\mathcal{D}^{T}=\{(x_{j}^{T},y_{j}^{A,T})\}_{j=1}^{m}$
  • Pre-trained MoE-LLM $\pi_{\theta}$ after modality alignment
2:Specialized expert allocation strategy $\{\mathbf{E}_{l}^{A},\mathbf{E}_{l}^{T}\}_{l=1}^{L}$ {$L$ is the total number of layers.}
3:Initialize: $k\leftarrow\text{\# audio experts}$, $M\leftarrow\text{\# total experts}$, $C\leftarrow\text{\# activated experts}$
4:Phase 1: Expert Selection Statistics
5:Extract hidden states: $\mathbf{h}_{l}^{A}=\pi_{\theta}(\mathcal{D}^{A})$, $\mathbf{h}_{l}^{T}=\pi_{\theta}(\mathcal{D}^{T})$ {Layer-wise representations}
[…]
7:Initialize counting matrices: $\mathbf{E}^{A}_{l}\leftarrow\mathbf{0}_{M}$, $\mathbf{E}^{T}_{l}\leftarrow\mathbf{0}_{M}$ {Per-layer expert counters}
8:for $(modality,\mathbf{h}_{l})$ in $\{(A,\mathbf{h}_{l}^{A}),(T,\mathbf{h}_{l}^{T})\}$ do
9:for $i\leftarrow 1\text{ to }T^{modality}$ do
10:  $\mathcal{E}_{l,i}^{modality}\leftarrow\mathrm{top\text{-}k}\big(\mathrm{MoE}(\mathbf{h}_{l}^{modality})_{i},C\big)$ {MoE using equation 3}
11:  for $j\leftarrow 1\text{ to }\text{len}(\mathcal{E}_{l,i}^{modality})$ do
12:   $\mathbf{E}^{modality}_{l}[\mathcal{E}_{l,i}^{modality}[j]]\mathrel{+}=1$ {Aggregate selections}
13:  end for
14:end for
15:end for
16:Phase 2: Modality Token Load Ratio Computation
17:for $modality\in\{A,T\}$ do
18:for $j\leftarrow 1\text{ to }M$ do
19:  $\rho_{l,j}^{modality}\leftarrow\frac{\mathbf{E}^{modality}_{l,j}}{C\cdot T^{modality}}$ {Token load ratio $\rho\in[0,1]$}
20:end for
21:end for
22:Phase 3: Partitioning Modality Experts Based on Modality Token Load
23:for $j\leftarrow 1\text{ to }M$ do
[…]
26:end for
27:return $\pi_{\theta}(\mathrm{Audio\_Experts}_{l},\mathrm{Text\_Experts}_{l})$
@@ 3.3 Training Strategy for Modality Experts
Modality Aligment During the modality alignment stage, audio ASR data is used to align the semantic spaces of speech and text. In this stage, the main focus is on training the Adapter that connects the Audio Encoder and the LLM. Through modality alignment training, it is ensured that speech can be understood by the LLM in the same way as the meaning expressed by text. The specific approach is to use a cross-entropy loss to align the audio and text modalities. Because the audio sequence for the same content is significantly longer than the text, in addition to downsampling the audio, we also use text padding tokens to align their lengths.
@@ 3.4 Audio Generation with Reinforcement Learning
To broaden the coverage of speakers and domains, it is inevitable that the pre-training data contains label noise and pronunciation errors, which can lead to hallucinations in the model. We observed that the quality of generated response audio is not consistent under different sampling parameters. To ensure more stable quality in response audio generation, we introduce a reinforcement learning stage to improve the stability of speech generation. Specifically, we use the text from the LibriSpeech Panayotov et al. (2015) set and have the model paraphrase it to obtain response audio and corresponding text. The response audio is then transcribed into recognized text using Whisper-large, and the word error rate (WER) is calculated as the reward score. All samples are ranked according to their WER, and triplet preference data $(x,y^{A}_{w},y_{l}^{A})\sim D$ is constructed for direct preference optimization (DPO) Rafailov et al. (2023) training as follows:
@@
where $\theta$ is the policy model and refref is the reference model. $x$ denotes the text input of the audio to be synthesized, $y_{w}^{A}$ denotes a good speech sequence, and $y_{l}^{A}$ denotes a bad speech sequence.
@@ 4.1 Experiment Settings
Models The overall architecture of the DeepOmni model is shown in Figure 2. We use the pre-trained DeepSeek-V2-Lite as the initial MoE-based LLM backbone, Whisper-medium as the audio encoder, and an MLP-based adapter to connect the audio encoder and the LLM. DeepOmni contains a total of 66 experts, including 6 audio experts, 58 text experts, and 2 shared experts. A sparse routing mechanism is adopted, where each token is processed by 6 experts. The model consists of 27 layers in total: the first layer is a dense layer, followed by 26 MoE layers. Each MoE layer has a hidden size of 1408, and the self-attention layers use flash-attention Dao (2023) with 16 attention heads. In stage 1, the learning rate is set to 2e-5. Since the original LLM has not undergone any pre-training on audio modalities, the learning rate is set to 1e-4 in stage 2.1 to accelerate convergence, 2e-5 in stage 2.2, and 5e-5 in stage 3. We use the AdamW Loshchilov & Hutter (2017) optimizer with $\beta_{1}$= 0.9 and $\beta_{2}$ = 0.95. The final evaluation is conducted on LLM benchmarks, SQA, ASR, and TTS tasks.
Datasets In stage 1, WenetSpeech Zhang et al. (2022) is used for modality alignment. In stage 2.1, AudioQA-1M Gao et al. (2025) is used for specialized training of audio experts. In stage 2.2, databricks-dolly-15k Dolly (2023), MathInstruct Yue et al. (2023) (262K) and camel-ai-math Li et al. (2023) (50K) are used for specialized training of text experts. In stage 3, AudioQA-1M is used for joint training of modality experts. For the RL stage, we sampled 28K text entries from LibriSpeech Panayotov et al. (2015) to construct the audio preference pair dataset.
@@ 4.2 Evaluations
Evaluation on Text to Text The text evaluation is mainly conducted to assess the retention of language capabilities in the models, and all evaluations are performed using opencompass [^1] https://opencompass.org.cn/. As shown in Table 3, modular MLLMs cause minimal damage to the language abilities of the original text LLMs, primarily because the LLM is only responsible for generating text tokens. However, native MLLMs significantly impair the language capabilities of the original text LLMs, with an average relative loss of over 20% compared to their backbone LLMs. The DeepOmni model effectively alleviates the catastrophic forgetting problem of native MLLMs through modality isolation and joint modality training, achieving an average relative loss at the same level as modular MLLMs.
@@
Model #A-Param LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Modular MLLMs Minmo Chen et al. (2025) 7B 78.9 64.1 55.0 39.9 48.3 37.5 60.7 47.2 Step-Audio Huang et al. (2025) 130B 81.0 - 75.1 - 58.0 - 71.4 - Freeze-Omni Wang et al. (2024) 7B 72.0 - 44.7 - 53.9 - 56.9 - Native MLLMs Moshi Défossez et al. (2024) 7B 62.3 21.0 26.6 9.2 22.8 7.3 37.2 12.5 GLM-4-Voice Zeng et al. (2024) 9B 64.7 50.7 32.2 15.9 39.1 26.5 45.3 31.0 LUCY Gao et al. (2025) 7B 59.6 51.0 26.6 18.2 23.2 18.2 36.5 29.1 MoExtend Zhong et al. (2024) 2.4B 54.0 45.3 20.8 16.9 18.5 17.2 31.1 26.5 DeepOmni 2.4B 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
The evaluation of speech-to-text is divided into two tasks: Spoken QA and ASR. As shown in the S $\rightarrow$ T column of Table 2, thanks to the preservation of the original LLM’s language capabilities by modular MLLMs, these models generally achieve leading performance. In contrast, native MLLMs cause significant degradation to the original LLM, resulting in increased hallucinations and thus lower scores. However, DeepOmni effectively mitigates the damage caused by native multimodality to the LLM, preserving the language capabilities of the LLM to the greatest extent. As a result, DeepOmni leads among native MLLMs. The ASR capabilities of the models are shown in Table 4, where DeepOmni, with a relatively small number of parameters, achieves performance comparable to other models with much larger parameter sizes.
@@
Model #A-Params WenetSpeech AIShell LibriSpeech test-net test-meeting test test-clean test-other Qwen2-Audio-base Chu et al. (2024) 7B 7.64 8.40 1.52 1.74 4.04 Baichuan-Audio-base Li et al. (2025a) 7B 10.13 13.28 1.93 3.02 6.04 VITA Fu et al. (2025) 7B 12.15 16.53 - 8.14 18.41 Step-Audio-chat Huang et al. (2025) 130B 9.47 10.83 2.14 3.19 10.67 Qwen2.5-Omni Xu et al. (2025) 7B 6.04 7.71 1.13 2.37 4.21 GLM-4-Voice Zeng et al. (2024) 9B - - 3.02 2.10 4.90 DeepOmni 2.4B 8.98 10.67 2.42 3.25 8.21
Evaluation on Text to Speech As shown in Table 1, the last three rows present the results for DeepOmni, where ”topk” denotes the sampling parameter for audio tokens. Since the speech dialogue data used by DeepOmni contains a relatively high proportion of Chinese, the model performs better on Chinese tasks but is somewhat lacking in English. After optimizing audio generation during response with DPO, the model may already perform well on simple tasks, as DPO increases the probability of the correct audio appearing in pass@1, thereby making high-quality audio generation more stable. However, for more difficult tasks (test-hard), the supervised fine-tuning (SFT) model itself may not generate the correct result within the entire search space, so applying DPO brings little to no benefit.
Evaluation on Speech to Speech The evaluation of speech-to-speech is shown in the S $\rightarrow$ S column of Table 2. MoExtend refers to the results obtained by extending the original Deepseek-V2-Lite model with additional experts, adding six audio experts. However, since expanding the number of experts alters the distribution of the router, the performance is relatively suboptimal. DeepOmni, on the other hand, adopts an adaptive modality expert selection strategy, preserving the original router’s allocation to experts. This allows the audio experts to specialize while retaining the capabilities of the original text experts, thus achieving better performance among native MLLMs.
Latency We deployed DeepOmni on a web server and measured end-to-end latency in half-duplex mode, where users manually control the timing of their speech input, without involving a Voice Activity Detection (VAD) module. The reported numbers are averaged over more than 10 samples. With a good network connection, the end-to-end latency of the entire system in half-duplex mode is 0.4362 seconds. The time cost for generating the first audio chunk is 0.3417 seconds. By subtracting the first chunk cost from the total latency, we estimate the network transmission cost to be 0.1062 seconds. The time cost for decoding a single token is 0.0205 seconds. All time measurements were conducted on a GPU with 80GB of memory.
@@ 5 Conclusion
To address the catastrophic forgetting problem caused by the scarcity of paired data, we propose DeepOmni, a MoE framework for adaptive modality expert learning. DeepOmni first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepOmni incurs a performance loss lower than peer native MLLMs and on par with modular ones. Meanwhile, the end-to-end dialogue latency remains within 0.5 seconds, ensuring a seamless and intelligent speech interaction experience. To the best of our knowledge, this is the first MoE-based native multimodal large speech interaction model, which opens up a new direction for mitigating catastrophic forgetting in native multimodal speech interaction models and bridges the gap between native multimodal and modular multimodal approaches.
6 Reproducibility statement
We have released all reproducible code in the supplementary materials and anonymized it. The details of the training data used and the hyperparameters are provided in Appendices A.1 and B, respectively.
Appendix A Appendix
A.1 Dataset Details
@@ Appendix B Experimental Setup Details
Table 6: Hyperparameters used in each stage training on DeepOmni.
@@
B.1 Ablation Study
B.1.1 Audio Experts Number
@@
B.1.2 The Modality Load of Different Layer Experts
Figure 4: The modality load of the experts in the first layers.
@@ B.1.2 The Modality Load of Different Layer Experts
The expert loads at different layers are shown in Figures  4,  5, and  6, corresponding to the modality loads of the first, middle, and last MoE layers, respectively. We can see that the modality loads differ substantially across layers, indicating that the indices of the audio experts vary from layer to layer. Moreover, the modality token loads at each layer also reveal the specialization of the modality-specific experts.
B.1.3 Modality-Specific MoE